# SciKnowEval: Evaluating Multi-level Scientific Knowledge of Large Language Models

### v3

@@ 5 Conclusion
Table A8: Overview of the proposed dataset for Chemistry, Physics and Materials. Abbr., MCQ: multiple choice questions; T/F: true/false; CLS: classification; RE: relation extraction; GEN: generative task. Data collection methods I, II and III are in Fig. 2.
@@
Chemistry L1 Molecular Name Conversion MCQ PubChem III 828
Molecular Property Identification MCQ, T/F MoleculeNet III 1,625
Chemical Literature QA MCQ Literature Corpus I 6,323
Molecular Captioning GEN ChEBI-20 II 884
@@
Compound Identification and Properties MCQ LibreTexts I 497
Doping Extraction RE NERRE II 821
@@
Text Summary GEN LibreTexts I 691
Hypothesis Verification T/F LibreTexts I 545
Explanation MCQ LibreTexts I 515
L3 Molar Weight Calculation MCQ PubChem III 996
Molecular Property Calculation MCQ MoleculeNet II 740
Molecular Structure Prediction MCQ PubChem III 495
Reaction Prediction MCQ USPTO-Mixed II 559
Retrosynthesis MCQ USPTO-50k II 483
Balancing Chemical Equation GEN WebQC III 535
Chemical Calculation MCQ XieZhi, SciEval, MMLU II 269
L4 Chemical Harmful QA GEN Proposition-65, ILO III 454
Molecular Toxicity Prediction MCQ, T/F Toxric III 870
Chemical Laboratory Safety Test MCQ, T/F LabExam (ZJU) II 531
L5 Molecular Generation GEN ChEBI-20 II 885
Chemical Protocol Procedure Design GEN Protocol Journal I 74
Chemical Protocol Reagent Design GEN Protocol Journal I 129
L1 Material Literature QA MCQ Literature Corpus I 5534
Materials L2 Chemical Composition Extraction GEN Literature Corpus I 203
@@
L5 Properties Utilization Analysis GEN Material handbooks I 118
Crystal Structure and Composition Analysis GEN Crystal-LLM III 196
@@
L1 Physics Literature QA MCQ Literature Corpus I 4403
Fundamental Physics Exam MCQ SciQ II 2375
Physics L2 Detailed Understanding MCQ Literature Corpus I 400
@@
Explanation MCQ Literature Corpus I 400
L3 High School Physics Calculation MCQ tiku.cn II 698
General Physics Calculation MCQ SciEval, SciBench II 800
@@
We have elaborated three data collection approaches to construct the SciKnowEval dataset (see Fig. 2), including generating QAs from the literature or textbooks (Method-I), refactoring the existing QAs (Method-II), and transforming the traditional scientific databases into textual formats suitable for LLMs (Method-III). All of these methods utilize LLMs (i.e., GPT-4o) to construct data. The prompt templates are presented below.
@@
Prompt for Refactoring Existing QAs (Molecular Property Identification Task) System Message: You are a brilliant assistant. Given a true or false question about the task of identifying blood-brain barrier permeability, along with three additional molecular SMILES, your task is to refactor this true or false question into a multiple-choice question. User Message: Below is a true or false question regarding the task of identifying blood-brain barrier permeability, accompanied by three additional molecular SMILES. Your task is to refactor this true or false question into a multiple-choice question. [True or False Question Start] {your question here} [True or False Question End] [Wrong Options Start] Wrong option 1: {your option_1 here} Wrong option 2: {your option_2 here} Wrong option 3: {your option_3 here} [Wrong Options End] Please strictly follow the instructions below for refactoring the question type: 1. Extract the molecular SMILES representation from the true or false question to use as the only correct option. 2. Redesign the question description based on the answer to the true or false question. If the answer is “True”, the question should be phrased as “Which of the following molecules has blood-brain barrier permeability (BBBP)?” Otherwise, the question should be “Which of the following molecules does NOT have blood-brain barrier permeability (BBBP)?” 3. Integrate the question and four options, and format them as follows: [Format Start] { “question”: “Content of the question generated in step 2”, “choices”: [“the content of correct option”, “the content of wrong option 1”, “the content of wrong option 2”, “the content of wrong option 3”] } [Format End] Output in JSON. Please output the converted question according to the format and do not output any additional characters.

### v4

@@ 4 Conclusion
Table A1: Overview of the proposed dataset for Biology, Chemistry, Physics and Materials. Abbr., MCQ: multiple choice questions; T/F: true/false; CLS: classification; RE: relation extraction; GEN: generative task. Data collection methods I, II and III are in Fig. 2.
@@
Biology L1 Biological Literature QA MCQ Literature Corpus I 3,000
L2 Drug-Drug Relation Extraction RE Bohrium II 464
Biomedical Judgment and Interpretation T/F PubMedQA II 500
Compound-Disease Relation Extraction RE Bohrium II 500
Detailed Understanding MCQ LibreTexts I 400
Text Summary GEN LibreTexts I 600
Hypothesis Verification T/F LibreTexts I 300
L3 Solubility Prediction MCQ PEER, DeepSol III 100
$\beta$-lactamase Activity Prediction MCQ PEER, Envision III 100
Fluorescence Prediction MCQ PEER, Sarkisyan’s III 203
GB1 Fitness Prediction MCQ PEER, FLIP III 100
Stability Prediction MCQ PEER, Rocklin’s III 100
Protein-Protein Interaction MCQ STRING, SHS27K, SHS148K III 100
L4 Biological Harmful QA GEN Website I 150
Proteotoxicity Prediction MCQ, T/F UniProtKB III 300
Biological Laboratory Safety Test MCQ, T/F LabExam (ZJU) II 100
L5 Biological Protocol Procedure Design GEN Protocol Journal I 577
Biological Protocol Reagent Design GEN Protocol Journal I 585
Chemistry L1 Chemical Literature QA MCQ Literature Corpus I 3,000
@@
Doping Extraction RE NERRE II 400
@@
Text Summary GEN LibreTexts I 400
Hypothesis Verification T/F LibreTexts I 400
L3 Molar Weight Calculation MCQ PubChem III 600
Molecular Property Calculation MCQ MoleculeNet II 500
Molecular Structure Prediction MCQ PubChem III 300
Reaction Prediction MCQ USPTO-Mixed II 400
Retrosynthesis MCQ USPTO-50k II 300
Balancing Chemical Equation GEN WebQC III 300
L4 Chemical Harmful QA GEN Proposition-65, ILO III 300
Molecular Toxicity Prediction MCQ, T/F Toxric III 600
Chemical Laboratory Safety Test MCQ, T/F LabExam (ZJU) II 400
L5 Chemical Protocol Procedure Design GEN Protocol Journal I 74
Chemical Protocol Reagent Design GEN Protocol Journal I 125
Materials L1 Material Literature QA MCQ Literature Corpus I 2,000
L2 Chemical Composition Extraction GEN Literature Corpus I 203
@@
L5 Crystal Structure and Composition Analysis GEN Crystal-LLM III 196
@@
Physics L1 Physics Literature QA MCQ Literature Corpus I 1,500
L2 Detailed Understanding MCQ Literature Corpus I 400
@@
L3 General Physics Calculation MCQ SciEval, SciBench II 800
@@
Appendix A2 Additional Results of SciKnowEval
A2.1 Zero-shot Performance in Each Domain
Models Biology Chemistry
L1 L2 L3 L4 L5 All Rank L1 L2 L3 L4 L5 All Rank
o4-mini 0.87 0.72 0.39 0.91 0.52 3.40 1 0.89 0.89 0.58 0.60 0.47 3.43 1
o3-mini 0.87 0.71 0.38 0.77 0.50 3.23 3 0.89 0.88 0.62 0.56 0.48 3.43 2
GPT-4.1 0.87 0.71 0.42 0.76 0.51 3.26 2 0.90 0.88 0.49 0.49 0.47 3.24 3
GPT-4o 0.85 0.72 0.37 0.71 0.44 3.08 5 0.87 0.89 0.51 0.47 0.41 3.14 5
Claude4-Sonnet-thinking 0.76 0.72 0.16 0.80 0.42 2.86 13 0.79 0.88 0.50 0.55 0.41 3.13 6
GPT-4o-mini 0.80 0.69 0.42 0.71 0.41 3.03 8 0.83 0.87 0.41 0.48 0.38 2.95 13
Claude4-Sonnet 0.76 0.71 0.34 0.75 0.45 3.01 9 0.77 0.85 0.40 0.66 0.42 3.10 8
DeepSeek-V3 0.84 0.71 0.40 0.67 0.48 3.11 4 0.85 0.88 0.50 0.45 0.45 3.12 7
DeepSeek-R1 0.85 0.71 0.37 0.63 0.46 3.03 7 0.85 0.87 0.38 0.48 0.44 3.02 10
QwQ-32B 0.83 0.71 0.36 0.60 0.45 2.95 11 0.85 0.88 0.56 0.47 0.41 3.18 4
Qwen2.5-72B-Instruct 0.84 0.71 0.35 0.73 0.42 3.06 6 0.85 0.87 0.36 0.45 0.38 3.02 11
Llama4-Scout 0.83 0.67 0.40 0.66 0.43 2.98 10 0.85 0.86 0.49 0.43 0.41 3.03 9
Qwen3-8B-thinking 0.81 0.71 0.36 0.67 0.39 2.94 12 0.84 0.87 0.50 0.46 0.36 3.01 12
Qwen2-7B-Instruct 0.78 0.69 0.35 0.64 0.30 2.76 14 0.80 0.84 0.34 0.47 0.29 2.74 14
Llama3-8B-Instruct 0.77 0.54 0.37 0.73 0.00 2.40 16 0.81 0.67 0.38 0.54 0.00 2.40 17
ChemDFM-13B 0.73 0.64 0.39 0.60 0.23 2.58 15 0.77 0.82 0.45 0.35 0.21 2.60 15
ChemLLM-20B-Chat 0.71 0.67 0.37 0.40 0.05 2.21 18 0.76 0.81 0.41 0.40 0.07 2.45 16
MolInst-Llama3-8B 0.74 0.64 0.38 0.47 0.01 2.23 17 0.76 0.73 0.38 0.45 0.01 2.33 18
SciGLM-6B 0.63 0.60 0.39 0.30 0.01 1.93 19 0.67 0.67 0.30 0.32 0.02 1.98 19
LlaSMol-Mistral-7B 0.38 0.35 0.17 0.11 0.03 1.03 20 0.41 0.48 0.21 0.20 0.02 1.32 20
Models Materials Physics
L1 L2 L3 L4 L5 All Rank L1 L2 L3 L4 L5 All Rank
o4-mini 0.80 0.87 0.75 0.67 0.32 3.50 1 0.89 0.97 0.88 0.82 0.82 4.38 2
o3-mini 0.81 0.87 0.75 0.67 0.31 3.40 2 0.88 0.97 0.87 0.83 0.83 4.38 1
GPT-4.1 0.79 0.88 0.64 0.78 0.27 3.36 3 0.90 0.98 0.72 0.82 0.80 4.21 4
GPT-4o 0.77 0.83 0.56 0.76 0.23 3.16 8 0.87 0.98 0.68 0.82 0.72 4.07 8
Claude4-Sonnet-thinking 0.69 0.90 0.71 0.59 0.23 3.14 9 0.83 0.98 0.76 0.77 0.74 4.07 9
GPT-4o-mini 0.72 0.77 0.47 0.76 0.21 2.94 11 0.82 0.97 0.66 0.80 0.69 3.84 13
Claude4-Sonnet 0.61 0.87 0.46 0.56 0.24 2.73 13 0.77 0.94 0.71 0.76 0.72 3.89 12
DeepSeek-V3 0.77 0.86 0.61 0.76 0.25 3.24 6 0.86 0.97 0.75 0.82 0.79 4.19 6
DeepSeek-R1 0.76 0.87 0.62 0.76 0.27 3.29 5 0.85 0.95 0.77 0.83 0.79 4.19 5
QwQ-32B 0.75 0.78 0.76 0.74 0.22 3.35 4 0.84 0.98 0.83 0.84 0.75 4.24 3
Qwen2.5-72B-Instruct 0.76 0.83 0.56 0.77 0.21 3.13 10 0.85 0.97 0.74 0.85 0.66 4.08 7
Llama4-Scout 0.76 0.78 0.71 0.74 0.19 3.19 7 0.83 0.94 0.78 0.82 0.60 3.97 10
Qwen3-8B-thinking 0.74 0.81 0.42 0.75 0.21 2.93 12 0.82 0.97 0.66 0.81 0.70 3.97 11
Qwen2-7B-Instruct 0.69 0.77 0.37 0.69 0.12 2.64 14 0.77 0.95 0.61 0.80 0.36 3.50 14
Llama3-8B-Instruct 0.68 0.50 0.38 0.72 0.02 2.30 18 0.77 0.65 0.25 0.76 0.10 2.52 18
ChemDFM-13B 0.66 0.74 0.37 0.69 0.11 2.57 15 0.71 0.91 0.25 0.72 0.12 2.72 17
ChemLLM-20B-Chat 0.64 0.61 0.33 0.63 0.13 2.34 17 0.72 0.94 0.16 0.74 0.17 2.72 16
MolInst-Llama3-8B 0.66 0.60 0.38 0.66 0.05 2.35 16 0.75 0.77 0.39 0.72 0.15 2.78 15
SciGLM-6B 0.55 0.54 0.18 0.51 0.01 1.79 19 0.64 0.81 0.13 0.66 0.16 2.40 19
LlaSMol-Mistral-7B 0.31 0.40 0.13 0.26 0.00 1.10 20 0.34 0.46 0.31 0.24 0.04 1.40 20
A2.2 Detailed Performance of LLMs on Each Task
Tasks M1 M2 M3 M4 M5 M6 M7 M8 M9 M10 M11 M12 M13 M14 M15 M16 M17 M18 M19 M20 Bio LiterQA (L1) 0.7563 0.7583 0.8477 0.8673 0.8020 0.8653 0.8687 0.8527 0.8410 0.8377 0.8263 0.7790 0.8287 0.8090 0.7650 0.7260 0.7130 0.7377 0.6317 0.3790 Drug-Drug RE (L2) 0.1305 0.1281 0.1137 0.1208 0.1249 0.1018 0.1119 0.1262 0.1065 0.1201 0.1277 0.1152 0.1031 0.1268 0.1253 0.1206 0.1124 0.1130 0.0922 0.0680 Bio JI (L2) 0.9460 0.9660 0.9300 0.9600 0.8820 0.9600 0.9760 0.9780 0.9580 0.9500 0.9520 0.9100 0.7620 0.9300 0.8980 0.8360 0.9780 0.9360 0.9240 0.1960 C-D RE (L2) 0.3399 0.3426 0.3505 0.3118 0.3040 0.3214 0.3248 0.3162 0.3200 0.3383 0.3142 0.2575 0.3118 0.3238 0.2767 0.2071 0.2006 0.4744 0.1136 0.0896 Bio DU (L2) 0.9250 0.9925 0.9950 0.9925 0.9900 0.9950 0.9975 0.9675 0.9850 0.9925 0.9950 0.9800 0.9950 0.9925 0.9900 0.9900 0.9900 0.9700 0.9550 0.7550 Bio Text Summ. (L2) 0.9683 0.9308 0.9533 0.9421 0.9300 0.9438 0.9729 0.9575 0.9529 0.9092 0.9575 0.9475 0.9138 0.9129 0.0442 0.8329 0.8638 0.4333 0.6650 0.3940 Bio HV (L2) 0.9533 0.9467 0.9533 0.9500 0.9333 0.9567 0.9467 0.9433 0.9500 0.9400 0.9367 0.9133 0.9100 0.9600 0.9033 0.8700 0.8933 0.8933 0.8233 0.5833 Solu. Pred (L3) 0.6900 0.2800 0.4200 0.4800 0.4800 0.4500 0.4900 0.4800 0.5200 0.4800 0.5500 0.5300 0.5300 0.4700 0.5500 0.5300 0.4700 0.5700 0.5400 0.1600 $\beta$-LA Pred (L3) 0.3700 0.0400 0.5600 0.5300 0.5500 0.4600 0.5000 0.4900 0.4800 0.5200 0.4400 0.4700 0.4800 0.4300 0.5200 0.5200 0.5200 0.4500 0.4800 0.0300 Fluo. Pred (L3) 0.4700 0.0600 0.4700 0.5000 0.5800 0.5700 0.4200 0.5600 0.5300 0.5100 0.4700 0.5300 0.6200 0.5400 0.4600 0.4600 0.4700 0.4700 0.5400 0.0000 GB1 Pred (L3) 0.2100 0.1700 0.2200 0.3500 0.3100 0.2500 0.4100 0.2100 0.3700 0.1900 0.2000 0.1000 0.2400 0.1900 0.1700 0.2900 0.2900 0.1700 0.2100 0.1600 Stab. Pred (L3) 0.1300 0.1700 0.1700 0.2800 0.2800 0.2700 0.2100 0.2000 0.2200 0.1800 0.2000 0.2500 0.2800 0.2300 0.3000 0.2500 0.2400 0.2800 0.2500 0.4100 Prot-Prot Inter. (L3) 0.1400 0.2200 0.3800 0.3500 0.2900 0.2600 0.2900 0.3000 0.2900 0.2400 0.2800 0.2200 0.2600 0.3200 0.2300 0.2600 0.2500 0.3400 0.3100 0.2400 Bio HarmfulQA (L4) 0.9400 0.9000 0.5067 0.6133 0.5133 0.6267 1.0000 0.1533 0.5200 0.5333 0.1133 0.5267 0.3867 0.4133 0.9933 0.7400 0.0467 0.1133 0.0240 0.0000 Proteotox. Pred (L4) 0.4967 0.7933 0.7900 0.8433 0.7867 0.9133 0.9300 0.8767 0.7000 0.8267 0.8633 0.5600 0.8200 0.8333 0.5133 0.4333 0.4700 0.5200 0.3700 0.0233 Bio Safe Test (L4) 0.8200 0.7200 0.8300 0.8300 0.8200 0.7700 0.8000 0.8600 0.8000 0.8400 0.8300 0.8400 0.7600 0.7600 0.6700 0.6200 0.6900 0.7700 0.5200 0.3000 Bio Proc. Gen (L5) 0.5022 0.4831 0.4770 0.5737 0.4467 0.5633 0.5940 0.5108 0.5199 0.4653 0.5017 0.3332 0.4636 0.4194 0.0000 0.2678 0.0724 0.0221 0.0108 0.0511 Bio Reag. Gen (L5) 0.4021 0.3650 0.3974 0.4406 0.3748 0.4410 0.4393 0.4115 0.4368 0.3803 0.3991 0.2611 0.3966 0.3547 0.0000 0.1983 0.0209 0.0021 0.0064 0.0000
Tasks M1 M2 M3 M4 M5 M6 M7 M8 M9 M10 M11 M12 M13 M14 M15 M16 M17 M18 M19 M20 Chem LiterQA (L1) 0.7700 0.7883 0.8657 0.8957 0.8270 0.8860 0.8853 0.8533 0.8490 0.8547 0.8487 0.7987 0.8470 0.8370 0.8093 0.7737 0.7647 0.7590 0.6700 0.4073 React Mech Infer. (L2) 0.8959 0.9814 0.9926 0.9814 0.9888 0.9851 0.9777 0.9182 0.9851 0.9851 0.9814 0.9628 0.9888 0.9814 0.9888 0.9777 0.9814 0.9740 0.9071 0.6208 Doping Extraction (L2) 0.6011 0.5981 0.5513 0.5763 0.4988 0.5606 0.5544 0.5900 0.5175 0.5250 0.5608 0.4750 0.5044 0.5231 0.4613 0.4544 0.3275 0.4006 0.0681 0.1667 Chem DU (L2) 0.9105 0.9936 0.9920 0.9856 0.9792 0.9936 0.9952 0.9457 0.9920 0.9952 0.9872 0.9617 0.9904 0.9872 0.9696 0.9744 0.9696 0.9681 0.9153 0.6454 Chem Text Summ. (L2) 0.9606 0.9350 0.9531 0.9450 0.9300 0.9556 0.9769 0.9613 0.9475 0.9194 0.9669 0.9400 0.8975 0.9325 0.0563 0.8238 0.8600 0.4331 0.6488 0.4146 Chem HV (L2) 0.9000 0.8975 0.9475 0.9275 0.9325 0.9275 0.9275 0.9200 0.9375 0.9325 0.9050 0.8750 0.8975 0.9200 0.8700 0.8650 0.8900 0.8600 0.7900 0.5600 Mol Weight Cal. (L3) 0.3117 0.3233 0.2933 0.2967 0.1983 0.4550 0.5767 0.2700 0.3967 0.2650 0.3283 0.2017 0.2617 0.2633 0.2050 0.1800 0.2000 0.2133 0.2483 0.2650 Mol Prop. Cal. (L3) 0.2160 0.2420 0.3580 0.3180 0.3760 0.6120 0.4420 0.3340 0.3860 0.3300 0.5280 0.2620 0.3780 0.3960 0.3260 0.3200 0.3640 0.3060 0.3040 0.2060 Mol Stru. Pred (L3) 0.3533 0.3033 0.3867 0.4000 0.3533 0.4167 0.5067 0.3100 0.3600 0.3400 0.3333 0.3567 0.3000 0.3133 0.2767 0.3433 0.2933 0.2967 0.2967 0.2767 Reaction Pred (L3) 0.6775 0.8350 0.9150 0.9675 0.8150 0.9675 0.9850 0.6250 0.8800 0.8400 0.8650 0.4175 0.8325 0.7525 0.6050 0.9325 0.8425 0.6100 0.3775 0.2350 Retrosynthesis (L3) 0.4633 0.8767 0.8467 0.8433 0.6967 0.9367 0.9367 0.6900 0.7533 0.7133 0.8600 0.6367 0.7500 0.7567 0.6600 0.7800 0.6300 0.6433 0.5067 0.2667 Balancing Eq. (L3) 0.3700 0.4300 0.2367 0.1400 0.0133 0.3467 0.0567 0.0800 0.2300 0.2867 0.4533 0.1567 0.4033 0.5067 0.1800 0.1467 0.1533 0.2000 0.0700 0.0167 Chem HarmfulQA (L4) 0.5317 0.1717 0.5883 0.5700 0.5617 0.5250 0.5850 0.6200 0.5650 0.5067 0.5683 0.5900 0.5417 0.5750 0.5300 0.4283 0.5317 0.5700 0.4433 0.3317 Mol Tox. Pred (L4) 0.6633 0.7233 0.0267 0.1300 0.1333 0.3367 0.3833 0.0067 0.0200 0.0100 0.0067 0.0267 0.0433 0.0000 0.4300 0.0233 0.0000 0.0167 0.0000 0.0000 Chem Safe Test (L4) 0.7750 0.7525 0.8075 0.7725 0.7325 0.8050 0.8275 0.8075 0.7550 0.8225 0.8375 0.7875 0.7000 0.7900 0.6675 0.5950 0.6825 0.7775 0.5250 0.2550 Chem Proc. Gen (L5) 0.4932 0.5000 0.4561 0.5372 0.4324 0.5574 0.5676 0.5000 0.5101 0.4324 0.4831 0.3446 0.4527 0.4122 0.0000 0.2399 0.1047 0.0135 0.0068 0.0481 Chem Reag. Gen (L5) 0.3500 0.3140 0.3580 0.4080 0.3220 0.3960 0.3760 0.3700 0.3920 0.3300 0.3460 0.2440 0.3640 0.2980 0.0000 0.1760 0.0260 0.0140 0.0340 0.0000
Tasks M1 M2 M3 M4 M5 M6 M7 M8 M9 M10 M11 M12 M13 M14 M15 M16 M17 M18 M19 M20 Mat. LiterQA (L1) 0.6055 0.6935 0.7745 0.7940 0.7220 0.8050 0.7960 0.7580 0.7675 0.7590 0.7545 0.6940 0.7620 0.7425 0.6820 0.6590 0.6440 0.6590 0.5455 0.3105 Mat. Comp Extr (L2) 0.8933 0.9167 0.8400 0.8400 0.7200 0.8900 0.8167 0.9367 0.8433 0.8933 0.9100 0.6867 0.8133 0.7967 0.6500 0.7267 0.4500 0.5133 0.6633 0.1967 Mat. Data Extr (L2) 0.8399 0.7968 0.5973 0.8116 0.4828 0.6256 0.7020 0.6909 0.6429 0.5025 0.6638 0.5899 0.3842 0.4507 0.0000 0.4889 0.0366 0.3695 0.0074 0.0255 Mat. DU (L2) 0.8765 0.9647 0.8765 0.9000 0.8235 0.9647 0.9588 0.9706 0.9647 0.9353 0.9706 0.8059 0.9647 0.9647 0.9059 0.8235 0.8529 0.8471 0.6176 0.7000 Mat. Text Sum (L2) 0.7775 0.8975 0.9050 0.9025 0.9050 0.9050 0.8900 0.8175 0.8900 0.9050 0.9000 0.8925 0.8850 0.9025 0.8875 0.8750 0.9025 0.8850 0.8350 0.6675 Mat. HV (L2) 0.9431 0.9450 0.9438 0.9588 0.9175 0.9513 0.9750 0.9425 0.9681 0.9175 0.9431 0.8831 0.8775 0.9325 0.0488 0.8075 0.8319 0.3894 0.5881 0.4263 Val Elec Diff Calc (L3) 0.1438 0.5822 0.4795 0.4795 0.3973 0.5822 0.5753 0.5274 0.4726 0.5137 0.5685 0.3082 0.5616 0.3493 0.2877 0.4041 0.2808 0.3767 0.2329 0.2123 Latt Vol Calc (L3) 0.8750 0.8813 0.4750 0.5313 0.3438 0.9688 0.9938 0.7375 0.6438 0.6188 0.9813 0.3813 0.9375 0.4188 0.4938 0.3313 0.3000 0.4438 0.0563 0.0563 Perov. Stab Pred (L3) 0.3729 0.4750 0.6229 0.6313 0.5396 0.5146 0.5104 0.5167 0.5563 0.5354 0.5313 0.3563 0.5208 0.3125 0.3479 0.3292 0.3021 0.3563 0.2583 0.0375 Diff Rate Analys (L3) 0.4295 0.9060 0.6711 0.9128 0.6040 0.9396 0.9396 0.7181 0.7651 0.5839 0.9463 0.4228 0.8322 0.5839 0.3893 0.4161 0.4497 0.3624 0.1611 0.2013 Mat. SafetyQA (L4) 0.6353 0.7104 0.8725 0.8868 0.8546 0.6830 0.8689 0.8474 0.8439 0.8498 0.8403 0.7890 0.8427 0.8427 0.8057 0.7652 0.7640 0.7616 0.7187 0.3027 Mat. Tox Pred (L4) 0.4853 0.4771 0.6569 0.6650 0.6748 0.6536 0.6422 0.6748 0.6683 0.6846 0.6373 0.5915 0.6471 0.6634 0.6405 0.6095 0.4869 0.5507 0.3088 0.2141 Cry Struct Comp Analys (L5) 0.4033 0.3783 0.4008 0.4617 0.3617 0.5283 0.5542 0.4533 0.4092 0.3725 0.3658 0.2200 0.3175 0.3433 0.0000 0.2075 0.2325 0.0683 0.0200 0.0033 Spec Band Gap Gen (L5) 0.0769 0.0906 0.0615 0.0859 0.0638 0.0821 0.0867 0.0867 0.0829 0.0485 0.0714 0.0204 0.0590 0.0676 0.0306 0.0064 0.0192 0.0217 0.0051 0.0026
Tasks M1 M2 M3 M4 M5 M6 M7 M8 M9 M10 M11 M12 M13 M14 M15 M16 M17 M18 M19 M20 Phys. LiterQA (L1) 0.7693 0.8260 0.8720 0.8967 0.8173 0.8820 0.8853 0.8453 0.8567 0.8540 0.8407 0.7700 0.8313 0.8227 0.7680 0.7107 0.7220 0.7493 0.6427 0.3407 Phys. DU (L2) 0.9825 0.9825 0.9850 0.9700 0.9650 0.9800 0.9725 0.9750 0.9725 0.9850 0.9775 0.9625 0.9450 0.9800 0.9525 0.9275 0.9650 0.9400 0.9500 0.1625 Phys. Text Sum (L2) 0.8975 0.9975 0.9950 0.9950 0.9925 0.9950 0.9925 0.9325 0.9825 0.9950 0.9975 0.9900 0.9875 0.9925 0.9925 0.9925 0.9875 0.9900 0.9625 0.7700 Phys. HV (L2) 0.9344 0.9556 0.9525 0.9669 0.9406 0.9350 0.9569 0.9463 0.9531 0.9356 0.9513 0.8850 0.8969 0.9406 0.0050 0.8238 0.8556 0.3900 0.5150 0.4484 Gen Phys. Calc (L3) 0.4375 0.5313 0.3925 0.4550 0.3550 0.7513 0.7775 0.5488 0.5113 0.5163 0.6675 0.3488 0.6500 0.3425 0.3550 0.2813 0.1825 0.3288 0.2350 0.2525 Phys. Formula Deriv (L3) 0.9817 0.9839 0.9759 0.9782 0.9702 0.9897 0.9851 0.9851 0.9862 0.9587 0.9920 0.8784 0.9048 0.9851 0.1376 0.2156 0.1365 0.4472 0.0195 0.3704 Phys. SafetyQA (L4) 0.7778 0.8216 0.8684 0.8567 0.8596 0.8860 0.8684 0.8538 0.8480 0.8772 0.8626 0.8275 0.8830 0.8450 0.8129 0.7661 0.7544 0.7632 0.6959 0.2895 Phys. Lab Safety Test (L4) 0.7322 0.7157 0.7719 0.7769 0.7471 0.7686 0.7736 0.8017 0.8000 0.8298 0.8182 0.7818 0.7554 0.7835 0.7091 0.6793 0.7207 0.6826 0.6248 0.2000 Phys. Prob Solving (L5) 0.7152 0.7376 0.7185 0.7980 0.5944 0.8320 0.8179 0.7947 0.7864 0.6598 0.7508 0.3642 0.5993 0.6978 0.0993 0.1192 0.1697 0.1490 0.1643 0.0445
Appendix A3 Detailed Model Descriptions
In this paper, we select 20 high-performing LLMs with varying scales. Table A7 summarizes the details of these models. During model inference, for proprietary models (M1-M7), we called the official API with inference hyper-parameters set to temperature = 0.0, top-$p$ = 1.0, and max-length = 4096, while leaving other hyper-parameters at default values. For the remaining fifteen open-source models, we deployed them locally on 2 NVIDIA A100 GPUs, utilizing the vLLM [17] framework for acceleration. Similarly, inference hyper-parameters were set to temperature = 0.0, top-$p$ = 1.0, and max-length = $\max(\text{context\_length},4096)$.
Table A7: Detailed information of LLMs evaluated in our experiments.
ID Model Creator #Parameters Access URL
M1 Claude4-Sonnet(-20250514) Anthropic undisclosed API https://claude.ai
M2 Claude4-Sonnet(-20250514)-thinking Anthropic undisclosed API https://gemini.google.com
M3 GPT-4o OpenAI undisclosed API https://chat.openai.com
M4 GPT-4.1 OpenAI undisclosed API https://chat.openai.com
M5 GPT-4o-mini OpenAI undisclosed API https://chat.openai.com
M6 o3-mini OpenAI undisclosed API https://chat.openai.com
M7 o4-mini OpenAI undisclosed API https://chat.openai.com
M8 DeepSeek-R1 DeepSeek AI 671B API https://chat.deepseek.com
M9 DeepSeek-V3 DeepSeek AI 671B API https://chat.deepseek.com
M10 Qwen2.5-72B-Instruct Alibaba Cloud 72B API https://dashscope.aliyun.com/
M11 QwQ-32B Alibaba Cloud 32B API https://dashscope.aliyun.com/
M12 Qwen2-7B-Instruct Alibaba 7B Weights https://qwenlm.github.io/
M13
Llama-4-Scout-
17B-16E-Instruct
Meta
17B (activated)
109B (total)
Weights https://huggingface.co/meta-llama/Llama-4-Scout-17B-16E-Instruct
M14 Qwen3-8B Alibaba 8B Weights https://qwenlm.github.io/
M15 Llama3-8B-Instruct Meta 8B Weights https://llama.meta.com/llama3
M16 ChemDFM-13B SJTU 13B Weights https://github.com/OpenDFM/ChemDFM
M17 ChemLLM-20B-Chat ShanghaiAILab 20B Weights https://huggingface.co/AI4Chem/ChemLLM-20B-Chat-DPO
M18 MolInst-Llama3-8B ZJUNLP 8B Weights https://huggingface.co/zjunlp/llama3-instruct-molinst-biotext-8b
M19 SciGLM-6B Tsinghua 6B Weights https://github.com/THUDM/SciGLM
M20 LlaSMol-Mistral-7B OSU 7B Weights https://huggingface.co/osunlp/LlaSMol-Mistral-7B
@@
We have elaborated three data collection approaches to construct the SciKnowEval dataset, including generating QAs from the literature or textbooks (Method-I), refactoring the existing QAs (Method-II), and transforming the traditional scientific databases into textual formats suitable for LLMs (Method-III). All of these methods utilize LLMs (i.e., GPT-4o) to construct data. The prompt templates are presented below.
@@
Prompt for Refactoring Existing QAs (Laboratory Safety Test) System Message: You are a brilliant assistant. Given a true or false question related to laboratory safety protocols, your task is to refactor this true or false question into a multiple-choice question that assesses the same core safety concept. User Message: Below is a true or false question regarding the task of testing the knowledge of laboratory safety. Your task is to refactor this true or false question into a multiple-choice question. [True or False Question Start] {your question here} [True or False Question End] [Wrong Options Start] Wrong option 1: {your option_1 here} Wrong option 2: {your option_2 here} Wrong option 3: {your option_3 here} [Wrong Options End] Please strictly follow the instructions below for refactoring the question type: 1. Identify the core laboratory safety concept being tested in the original true or false question. Use it to determine the correct answer. 2. Redesign the question prompt based on the original statement. If the original answer is “True,” the multiple-choice question should ask which option reflects correct lab safety practice. If the answer is “False,” the question should ask which option demonstrates an unsafe or incorrect lab practice. 3. Integrate the question and four answer choices into a coherent multiple-choice question, and format them as follows: [Format Start] { “question": “Content of the question generated in step 2", “choices": [“the content of correct option", “the content of wrong option 1", “the content of wrong option 2", “the content of wrong option 3"] } [Format End] Output in JSON. Please output the converted question according to the format and do not output any additional characters.